OpenStudy (dancergirl45):

Is the simplified form of 2 square root of 3 ⋅ square root of 12 rational?

OpenStudy (dancergirl45):

@math&ing001

OpenStudy (math&ing001):

\[2\sqrt{3}\sqrt{12}=2\sqrt{3}\sqrt{3*4}=2(\sqrt{3})^{2}\sqrt{4}\] What do you think ?

OpenStudy (dancergirl45):

yes

OpenStudy (math&ing001):

Correct !

OpenStudy (dancergirl45):

Thank you

OpenStudy (math&ing001):

Welcome !